### v1

# Keyed Chaotic Tensor Transformations for Secure and Attributable Neural Inference

Abstract
This work introduces a novel framework for secure and privacy-preserving neural network inference based on keyed chaotic dynamical transformations. The proposed method applies a deterministic, cryptographically seeded chaotic system to tensors, producing non-invertible, user-specific transformations that enable authenticated inference, tensor-level watermarking, and data attribution. This framework offers a scalable and lightweight alternative to conventional cryptographic techniques, and establishes a new direction for tensor-level security in AI systems.
Note: This is a preliminary preprint released to share the core ideas and invite feedback. Empirical validation and formal security analysis are planned for revisions.
1 Introduction
Modern machine learning systems increasingly operate in settings where user privacy, data security, and attribution are critical. inference pipelines often transmit sensitive user data to hosted models. In these contexts, users must trust that their data are handled securely, and that their contributions are not exposed, misused, or misattributed. Existing approaches to privacy-preserving inference often rely on heavyweight cryptographic methods such as homomorphic encryption [1], and secure multi-party computation [2], or introduce performance trade-offs through techniques like differential privacy [3] and training-time watermarking [4]. These approaches can require significant computational overhead, retraining, or infrastructure changes, limiting their practicality in real-world inference pipelines.
To the best of my knowledge, no existing method offers a lightweight, deterministic approach to tensor-level security that supports both input privacy and contribution traceability, while remaining fully compatible with standard model architectures. Ideally, such a method would act as a cryptographic wrapper for input embeddings, applied either prior to or after inference, and would support a wide range of downstream use cases including authentication, watermarking, attribution, and data contribution auditing.
This work introduces a framework for applying keyed chaotic dynamical transformations to tensors used in machine learning pipelines, enabling privacy-preserving, traceable, and non-invertible manipulation of data representations. Each transformation is derived from a cryptographically seeded dynamical system, represented as a computation graph that operates deterministically on the input. The transformation can be configured to either preserve task-relevant structure or intentionally obscure it, depending on the application, resulting in representations that are non-invertible, watermarked, and cryptographically bound to a key. These transformations can be applied generally to tensors and are designed to integrate cleanly with existing models, without modifying architecture or requiring retraining.
The main contributions of this work are:
  • A method for applying deterministic chaotic transformations to tensors using keyed graph dynamical systems
  • A demonstration of how these transformations enable authenticated inference, watermarking, and attribution
  • A discussion of downstream applications including secure data contribution, dynamic ML firewalls, and tensor hash functions
This framework bridges ideas from cryptography, dynamical systems, and representation learning, offering a new lens on how models and data can interact securely. Unlike traditional cryptographic methods, this approach is architecture-agnostic, low-overhead, and enables per-user control over inference contributions. The following sections outline the system design, present empirical results on secure inference and watermark robustness, and explore future directions in secure, chaos-based ML infrastructure.
2 Background and Related Work
2.1 Cryptography
Modern cryptographic systems rely on mathematical transformations that are deterministic, non-invertible, and key-dependent. These properties are used to ensure data confidentiality, integrity, and authenticity across a wide range of communication protocols [5]. While most cryptographic methods operate on discrete data (e.g., byte strings), the key ideas — particularly keyed deterministic transforms and collision-resistant signatures — are relevant when designing mechanisms for secure computation in neural systems.
Two foundational building blocks in this space are symmetric-key encryption and cryptographic hash functions. Symmetric schemes such as AES and HMAC rely on a shared secret key to encrypt or authenticate data, ensuring that only entities with access to the key can reproduce or verify the transformation [6, 7]. These methods are computationally efficient and deterministic, but are designed for bitwise or block-level data and do not preserve semantic or geometric structure — making them incompatible with neural network pipelines where data often lives in high-dimensional continuous tensor spaces.
Cryptographic hash functions such as SHA-256 offer another relevant abstraction: they produce outputs that are deterministic, sensitive to small input changes, and hard to invert or collide [8]. However, they map inputs to fixed-length binary strings and do not support inverse gradients, interpolation, or alignment with downstream learning tasks.
The method introduced in this work draws inspiration from these primitives but is fundamentally distinct: it operates in continuous space, supports tensor-valued inputs, and is task-preserving by design. By using keyed chaotic systems to transform tensors, the method introduces cryptographically inspired properties—such as key-dependent transformation and input sensitivity—while remaining lightweight and compatible with standard model architectures.
2.2 Privacy-Preserving Inference
As machine learning systems become increasingly deployed in real-world, cloud-based, and client-server settings, privacy-preserving inference has emerged as a critical area of research. In many applications, users transmit sensitive data — such as biometrics, medical signals, personal text, or proprietary feature representations — to centralized models for inference. This has raised concerns not only about data exposure, but also about the reusability, traceability, and control of these contributions once they enter the inference pipeline.
A number of approaches have been proposed to mitigate these risks, but each carries substantial trade-offs in performance, infrastructure complexity, or applicability. Techniques such as homomorphic encryption (HE) [1] allow models to perform inference directly on encrypted data, preserving confidentiality without revealing inputs to the model host. However, HE-based approaches are often computationally intensive and have orders of magnitude slower model inference, limiting their deployment to narrow use cases. Similarly, secure multi-party computation (SMPC) distributes computation across parties such that no individual party has full access to the data, but these methods typically require multiple rounds of communication, synchronized computation, and significant coordination overhead [2].
Other approaches focus on statistical privacy. For example, differential privacy (DP) introduces controlled noise into model training or inference pipelines to mask individual contributions [3]. While DP offers formal privacy guarantees, it is stochastic, lossy, and often difficult to apply meaningfully in one-off inference scenarios. Lastly, training-time watermarking techniques embed identifiable signatures into the model’s parameters during training, enabling ownership verification or post hoc misuse detection [4]. However, these methods are typically static, global, and tied to the model itself—making them difficult to extend to per-user or inference-time watermarking, and often requiring access to the model weights.
The approach introduced in this work departs from these paradigms. Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior to or following inference. These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally obfuscated, and task-compatible. While the transformations are reversible by the key holder (e.g., via subtraction or reapplication), they remain opaque to unauthorized parties without knowledge of the underlying system. This enables lightweight, key-authenticated inference, where only correctly transformed inputs yield meaningful results, and offers tensor-level control over privacy, traceability, and secure data contribution—without requiring architectural changes or incurring the computational costs of conventional cryptographic methods.
2.3 Chaotic Dynamical System
Chaotic dynamical systems are deterministic processes that exhibit sensitive dependence on initial conditions, nonlinear evolution, and long-term unpredictability despite being fully governed by known rules. They are widely studied in physics, biology, and nonlinear systems theory, and have seen limited but notable use in machine learning—primarily as components in reservoir computing.
In the context of this work, chaos is used as a security primitive rather than a model architecture. Chaos is associated with the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small changes in initial conditions. The key properties of chaotic systems that make them suitable for this purpose include:
  • Determinism: a fixed chaotic system produces consistent trajectories from identical inputs.
  • Sensitivity to initial conditions: ensures that different inputs (or different users) produce non-overlapping, difficult-to-invert trajectories.
  • Structural complexity: allows for transformation of input tensors in ways that preserve local information but obscure global reversibility.
This work draws inspiration from systems such as coupled map lattices, chaotic attractors, and graph-structured reservoirs, using them to implement transformations that are key-conditioned, lightweight, and difficult to approximate or reverse without full knowledge of the graph topology and dynamical rules. By applying these systems to input or output tensors, it becomes possible to encode user-specific behavior, embed cryptographic structure into data, and control the visibility or traceability of representations within inference pipelines.
3 Keyed Chaotic Transformation Construction
The goal of this method is to construct a graph-based dynamical system whose behavior is key-conditioned and exhibits characteristics of chaos—including sensitivity to initial conditions, deterministic unpredictability, and structural complexity. This system serves as a transformation mechanism for input or output tensors, enabling secure and traceable inference without modifying the model architecture. The construction process consists of sampling a graph topology, selecting a dynamical update rule, evaluating chaotic behavior (e.g., via Lyapunov exponents), and deterministically assembling the system using a cryptographic key. The following subsections outline each stage in detail.
3.1 Graph Topology Sampling
The chaotic system used for transformation is implemented over a graph, where nodes represent internal state variables and edges define the pattern of dynamical interaction. The topology of this graph is sampled deterministically from a cryptographic key, which seeds a pseudorandom number generator. The goal is to produce a fixed but key-specific structure that exhibits rich dynamical behavior when coupled with nonlinear update rules.
Given a cryptographic key $k\in\{0,1\}^{n}$, a pseudorandom seed is derived as
$\text{seed}=\texttt{Hash}(k)\bmod 2^{32},$ (1)
where Hash is a standard cryptographic hash function (e.g., SHA-256), and the result is used to initialize a pseudorandom number generator. This generator is then used to deterministically sample a graph $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
In this work, topologies are sampled from families such as Erdős–Rényi graphs and Watts–Strogatz small-world graphs, though the framework generalizes to other constructions. For Erdős–Rényi graphs, edges are sampled independently:
$\mathbb{P}\big{(}(i,j)\in E\big{)}=p\quad\text{for all }i\neq j.$ (2)
for a fixed edge probability $p\in(0,1)$ and edge list $E$. Optionally, edge weights $w_{ij}\in[-a,a]$ may be sampled and scaled to control spectral properties. For small-world graphs, a Watts–Strogatz procedure is used, beginning from a $k$-regular ring lattice with each edge rewired independently with probability $\beta$.
Graphs are directed, may be weighted or unweighted, and typically disallow self-loops. The number of nodes is chosen to match the dimensionality of the input tensor segment being transformed (e.g., embedding size or output dimensionality).
Preliminary observations suggest that the structure of the graph strongly influences the resulting dynamics—including stability, memory, and sensitivity to perturbation. For example, graphs with higher connectivity (or spectral radius) may produce more stable or diffusive behavior, while sparse or modular graphs may exhibit more chaotic evolution. However, a full characterization of the relationship between graph topology and dynamical behavior is left as future work and presents a rich area for both empirical and theoretical exploration.
3.2 Dynamical Rule Selection
Each node in the system evolves according to a discrete-time update rule that combines a local nonlinear map with global coupling from neighboring nodes. The resulting dynamical behavior emerges from the interaction between the node-level nonlinearity and the graph structure. After this update, an optional leaky integration step is applied to blend the new state with the previous one.
Let $\mathbf{x}(t)\in\mathbb{R}^{d}$ represent the state of the system at time $t$, and let $A\in\{0,1\}^{d\times d}$ and $W\in\mathbb{R}^{d\times d}$ denote the binary adjacency matrix and edge weight matrix, respectively. The state update proceeds in two stages:
3.2.1 Model-specific update rule:
$\tilde{\mathbf{x}}(t+1)=f\left(\mathbf{x}(t),A,W;\theta\right)$
where $f$ denotes the chosen node-level update rule (e.g., logistic map, Izhikevich, linear), and $\theta$ represents optional rule-specific parameters (e.g., $r$, $\gamma$, thresholds). In most cases, $f$ incorporates both local nonlinear updates and graph-mediated coupling via the weighted adjacency matrix.
3.2.2 Leaky integration (optional):
$\mathbf{x}(t+1)=(1-\alpha)\cdot\mathbf{x}(t)+\alpha\cdot\tilde{\mathbf{x}}(t+1)$
where $\alpha\in[0,1]$ is the leak rate.
In this work, several node-level maps are supported:
  • Logistic map: $f(x_{i})=rx_{i}(1-x_{i})$, where $r\in(3.57,4)$ controls chaotic intensity.
  • Linear map: $f(x_{i})=x_{i}$, for baseline comparison.
  • Izhikevich dynamics: A two-variable spiking neuron model used to explore biologically inspired dynamics.
At runtime, the update rule is selected from a library of registered dynamics, and its parameters are sampled from known regimes that exhibit chaotic behavior. This selection is deterministically conditioned on the cryptographic key, enabling the transformation system to be both modular and reproducible across users and devices.
3.3 Chaos Diagnostics and Characterization
To ensure that the graph dynamical systems constructed using cryptographic keys exhibit chaotic behavior, we perform a series of diagnostics drawn from nonlinear systems theory. Chaos is typically defined by the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small differences in initial conditions. This property is essential for ensuring that similar inputs or neighboring keys result in divergent transformation trajectories, providing both privacy and resistance to inversion.
3.3.1 Lyapunov Exponent Estimation.
To assess whether the system exhibits chaotic behavior, we estimate the largest Lyapunov exponent using a standard two-trajectory method. We initialize two states: $\mathbf{x}_{0}$ and a perturbed version $\hat{\mathbf{x}}_{0}=\mathbf{x}_{0}+\epsilon$, where $\epsilon$ is a small positive scalar. Both trajectories are evolved synchronously under the same dynamical system. At each step, the separation $\delta(t)=\hat{\mathbf{x}}(t)-\mathbf{x}(t)$ is measured, and the Lyapunov exponent is estimated by averaging the log growth rate:
$\lambda\approx\frac{1}{T}\sum_{t=1}^{T}\log\left(\frac{\|\delta(t)\|}{\epsilon}\right)$
To ensure numerical stability, the perturbed state is renormalized after each step:
$\hat{\mathbf{x}}(t)\leftarrow\mathbf{x}(t)+\epsilon\cdot\frac{\delta(t)}{\|\delta(t)\|+\varepsilon}$
where $\varepsilon$ is a small constant to avoid division by zero. A consistently positive value of $\lambda$ indicates exponential divergence and confirms the presence of chaotic dynamics.
3.4 Key-Conditioned System Assembly
The final stage of the construction process assembles the graph dynamical system by combining the sampled topology, selected dynamical rule, and initial state—all deterministically derived from a shared cryptographic key. This key serves as the sole source of entropy for the transformation and ensures that the system is uniquely bound to the key holder.
To maintain predictable chaotic behavior, the method restricts sampling to a curated set of parameter ranges known to yield high sensitivity and non-periodicity. Each candidate configuration includes:
  • A graph family (e.g., Erdős–Rényi, Watts–Strogatz)
  • Associated hyperparameters (e.g., edge probability $p$, rewiring rate $\beta$, node degree $k$)
  • A dynamical rule (e.g., logistic map, Izhikevich system)
  • Rule-specific parameters (e.g., $r$ for logistic map, coupling strength $\gamma$, integration constants)
These configurations are indexed, and a pseudorandom generator seeded from the key selects from this indexed set. This ensures that the resulting system is both deterministic and cryptographically unique, while also avoiding regimes where dynamics are trivial or periodic.
In addition to the structural parameters, the initial state of the system is randomly sampled using the same seed derived from the key. This means both the system’s structure and starting conditions are key-dependent, ensuring that even if two users select the same dynamical configuration, their trajectories will diverge due to differing initial conditions.
The result is a fully assembled, chaotic transformation system that is:
  • Deterministic: reproducible with the key, without requiring key storage
  • Secure: opaque to observers without the key
  • Flexible: compatible with a range of graph and dynamic structures
This construction enables lightweight, cryptographically bound transformations.
3.5 Security Properties
The proposed system provides the following properties:
  • Key-conditioned determinism: Each user’s transformation system is deterministically derived from their cryptographic key, ensuring reproducibility without shared state or explicit key transmission.
  • Input obfuscation: Transformed inputs are non-trivial to interpret without the key, due to both structural scrambling and sensitivity to initialization. Small changes in the key or input yield divergent outputs.
  • Attribution and binding: Only a key holder can produce a given transformation. This allows inference outputs to be linked to specific users or devices, supporting attribution and misuse detection.
  • Opaque trajectories: Without knowledge of the graph topology, weights, and update rule, reverse-engineering the transformation is computationally impractical. The chaotic dynamics ensure exponential divergence from any incorrect reconstruction attempt.
4 Applications
4.1 Key-Authenticated Neural Network Inference
In cloud-hosted or API-based inference systems, it is often desirable to restrict access to only authorized users—both for security and for usage control. Traditional access control mechanisms typically operate at the network or user level—using API keys, authentication tokens, or IP-based restrictions. While these methods can identify the requester, they provide no cryptographic binding between the user and the data submitted to the model. As a result, once access is granted, any input can be submitted, including replayed, forged, or malicious content.
The proposed transformation scheme shifts authentication to the data level by requiring that inputs be correctly transformed using a user-specific, key-conditioned chaotic system. This approach ensures that only users with the correct cryptographic key can produce inputs that yield valid model outputs, effectively binding model access to both the identity and integrity of the data itself. It offers an additional layer of protection against misuse, credential leakage, and unauthorized inference—without requiring changes to the model or encrypted communication channels.
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a sequence of token embeddings. A user constructs a chaotic transformation system from a cryptographic key $k$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$. The input is scrambled client-side via:
$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
and transmitted to the model. At inference time, the model internally reconstructs the true input by regenerating the same chaotic system and subtracting the mask:
$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
Only if the correct key is used will this inversion yield a valid input. Any incorrectly scrambled input results in faulty or degraded inference, effectively binding model access to possession of the correct key.
Optionally, the model can also apply a forward scrambling to its outputs using a second chaotic trajectory:
$\tilde{\mathbf{y}}=\mathbf{y}+\mathbf{S}^{\prime}_{k}$
The client then performs the inverse:
$\mathbf{y}=\tilde{\mathbf{y}}-\mathbf{S}^{\prime}_{k}$
This creates a fully authenticated inference pipeline, where both input and output representations are cryptographically tied to the user’s key. This transformation operates independently of network-layer encryption, and can be combined with standard secure transport protocols to protect data in transit.
4.2 Watermarking for Model & User Attribution
In multi-user inference systems, it is often desirable to verify which user or device generated a particular output—especially in contexts involving auditing, content moderation, or abuse prevention. While some systems incorporate output watermarking to identify model-generated content, these are typically model-wide and do not support per-user attribution.
The proposed framework enables per-user output watermarking by applying a deterministic, key-conditioned transformation to the model output. Let $\mathbf{y}\in\mathbb{R}^{d}$ represent the output of a model (e.g., a logit vector, embedding, or generated content representation). A user-specific watermark is applied as:
$\tilde{\mathbf{y}}=\mathbf{y}+\mathbf{S}^{\prime}_{k}$
where $\mathbf{S}^{\prime}_{k}\in\mathbb{R}^{d}$ is a deterministic mask generated from the user’s key using the same chaotic transformation framework. This watermark is embedded server-side before the output is returned to the user or downstream system.
Crucially, this transformation acts as a soft fingerprint: the output remains usable for its original task, but contains a key-specific signal that can later be used to attribute the output to a specific user. If a model output appears in the wild, the server or auditor can regenerate $\mathbf{S}^{\prime}_{k}$ for known keys and match it against the suspicious output to infer its origin.
This watermarking mechanism enables per-user attribution without modifying the model, embedding user IDs, or relying on metadata. It offers a lightweight path to traceable outputs in shared inference environments, content moderation systems, or multi-tenant deployment pipelines.
4.3 Data Privacy
Many inference systems operate on embeddings derived from sensitive user data, such as personal text, biometric signals, or proprietary representations. In these settings, raw embeddings are often transmitted to centralized models for downstream inference—posing a significant privacy risk if intercepted or logged. While full encryption can provide protection, it typically requires secure execution environments or homomorphic computation, which introduces significant overhead.
The proposed transformation scheme offers a lightweight alternative by obfuscating embeddings directly in their native vector space. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of token or patch embeddings. The user applies a chaotic, key-conditioned transformation as follows:
$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
where $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ is a deterministic mask generated from a graph dynamical system seeded by the cryptographic key $k$. The scrambled embeddings $\tilde{\mathbf{X}}$ retain the structure and dimensions required for downstream tasks, but are semantically obfuscated and difficult to interpret without knowledge of the system.
Importantly, because the transformation is deterministic and reversible with the key, the model hosted on the server can decode the input by regenerating the same mask internally:
$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
This obfuscation protects the user’s input from inspection in transit or at rest, while still enabling standard inference with minimal computational overhead. Furthermore, reverse engineering the obscured tensors is impractical without access to the key generating the chaotic transformation of the tensor hence securing the users raw data.
4.4 Secure Data Contribution to Shared Models
In collaborative machine learning systems—such as federated learning, crowd-sourced training, or multi-institutional data sharing—users may wish to contribute data to a central model while retaining control over their raw inputs. However, direct data sharing raises concerns about privacy, misattribution, and unauthorized reuse, particularly in domains involving personal or proprietary information.
The proposed transformation framework allows users to submit data obfuscated by a key-conditioned chaotic transformation. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of input embeddings, and let $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ be the mask generated from the user’s key $k$. The transformed data is computed as:
$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
and contributed to a shared model or data repository. The transformation preserves structure and dimensionality, but obfuscates semantic content, thereby protecting raw data from inspection or unauthorized reuse.
In some scenarios, the server may invert the transformation:
$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
to recover the original data for training. Alternatively, the model could be trained directly on $\tilde{\mathbf{X}}$, treating the transformation as a privacy-preserving encoding layer. However, training directly on obfuscated data may degrade performance depending on the task, and its effectiveness is likely to be application-specific. Further empirical study is needed to determine the conditions under which learning from transformed data remains effective.
5 Discussion
This work introduces a new class of tensor-level transformations based on keyed chaotic dynamical systems, designed to provide input obfuscation, output watermarking, and authenticated inference in modern machine learning pipelines. The system is lightweight, deterministic, and modular—requiring no changes to model architecture.
While the system provides functional security properties such as key-conditioned masking, user attribution, and trajectory-based unpredictability, it does not yet offer formal cryptographic guarantees. In particular, the transformation is reversible by design for authorized users, and may be vulnerable to chosen-input or model inversion attacks in some threat models. As such, it is best viewed as a complementary privacy mechanism—sitting between full encryption and raw data sharing—and may be combined with other techniques (e.g., differential privacy, TEEs, encrypted transport) in high-risk applications.
The choice of dynamical system (e.g., graph topology, update rule, initialization) plays a significant role in determining the transformation’s properties. While this work restricts itself to a curated set of configurations with empirically chaotic behavior, future work could explore learned dynamics, adversarial training of transformation functions, or differentiable chaos-conditioned layers for integration into model backbones.
More broadly, this work suggests that chaotic systems—often studied in the context of memory, control, or biological modeling—can be repurposed as lightweight cryptographic primitives for AI deployment. As the interface between private data and hosted models becomes increasingly important, new methods that offer user-level control, attribution, and secure signaling will be essential. This framework provides one such direction and opens several avenues for further development.
6 Conclusion
This work introduces a novel framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning. By leveraging deterministic chaos seeded from cryptographic keys, the method enables input obfuscation, output watermarking, and authenticated inference without modifying the underlying model. The approach is lightweight, task-compatible, and suitable for a range of real-world applications where privacy, attribution, or secure contribution is essential. As the need for secure AI deployment grows, this method offers a new building block at the interface of dynamical systems, cryptography, and neural computation.
7 Acknowledgements
The author gratefully acknowledges the UKRI PhD studentship at the University of Edinburgh, which supported prior research training and early exploration of related ideas. This work was completed independently during a period of academic leave. The author also acknowledges the use of large language models (LLMs) to support drafting, editing, and idea development during the research and writing process.